# Asymptotically normal estimators in high-dimensional linear regression

### v2

@@ 1 Introduction
Throughout the paper, we use the following notation. We write $k]=\{1,\ldots,k\}$ for $k\in\mathbb{N}$. For $q\in\mathbb{N}\cup\{\infty\}$, write $\|\bm{v}\|_{q}$ for the $\ell^{q}$ norm: for $\bm{v}=(v_{1},\ldots,v_{p})^{\top}\in\mathbb{R}^{p}$ with $p\in\mathbb{N}$, $\|\bm{v}\|_{q}=(\sum_{j=1}^{p}|v_{j}|^{q})^{1/q}$ $(q<\infty)$, $\|\bm{v}\|_{\infty}=\max_{1\leq j\leq p}|v_{j}|$. For an infinite sequence $\bm{v}=(v_{j})_{j\geq 1}\in\ell^{q}$, the same formulae apply. For a bounded linear operator $\bm{A}:\ell^{2}\to\ell^{2}$ or a matrix $\bm{A}$ in finite dimensions, let $\|\bm{A}\|_{\mathrm{op}}=\sup_{\|\bm{x}\|_{2}=1}\|\bm{A}\bm{x}\|_{2}$ denote the $\ell^{2}$‑induced operator norm. For a vector $\bm{v}=(v_{j})_{j\geq 1}$ and an index set $T\subset\mathbb{N}$ with $|T|<\infty$, write $\bm{v}_{T}=(v_{j})_{j\in T}$, which is viewed as an element of $\mathbb{R}^{|T|}$. In the finite‑dimensional case $\bm{v}\in\mathbb{R}^{p}$, we restrict to $T\subset[p]$. Similarly, for a $p\times p$ matrix $\bm{A}=(A_{ij})_{i,j\in[p]}$ and index sets $T,T^{\prime}\subset[p]$, write $\bm{A}_{T,T^{\prime}}=(A_{ij})_{i\in T,j\in T^{\prime}}$. For a random variable $X$, we write the sub-Gaussian norm of $X$, if it exists, as $\|X\|_{\psi_{2}}$, that is, $\|X\|_{\psi_{2}}=\inf\{C>0:\,\mathsf{E}[\exp({X^{2}}/{C^{2}})]\leq 2\}$. For a $d$-dimensional random vector $\bm{X}$, the sub-Gaussian norm of $\bm{X}$ is $\|\bm{X}\|_{\psi_{2}}=\sup_{\|\bm{u}\|_{2}=1}\|\langle\bm{u},\bm{X}\rangle\|_{\psi_{2}}$. For $d\in\mathbb{N}$, $\bm{\mu}\in\mathbb{R}^{d}$, and $\bm{V}\in\mathbb{R}^{d\times d}$ positive semi‑definite, we write $\mathcal{N}_{d}(\bm{\mu},\bm{V})$ for the $d$-dimensional normal distribution with mean $\bm{\mu}$ and covariance $\bm{V}$, with the subscript omitted when $d=1$. For random elements $X$ and $Y$ and a sequence $\{X_{n}\}_{n\in\mathbb{N}}$ of random elements, we write $X\stackrel{{\scriptstyle d}}{{=}}Y$ if $X$ and $Y$ have identical distributions, and we write $X_{n}\to^{d}X$ and $X_{n}\to^{p}X$ for convergence in distribution and in probability as $n\to\infty$, respectively.
@@ 2 Model set‑up and main result
where $\bm{\theta}_{n}\in\mathbb{R}^{p_{n}}$ and $\bm{Z}_{i,n}\in\mathbb{R}^{p_{n}}$ for $i\in[n]$. Assume that $\{(\bm{Z}_{i,n},\epsilon_{i,n})\}_{i\in[n]}$ are i.i.d. with $\mathsf{E}(\bm{Z}_{1,n})=\bm{0}$, $\mathsf{E}(\epsilon_{1,n})=0$, $\mathsf{V}(\epsilon_{1,n})=\sigma_{n}^{2}>0$, and $\bm{Z}_{1,n}$ independent of $\epsilon_{1,n}$. Let $\bm{\theta}_{0,n}$ denote the true parameter vector. Unless otherwise stated, throughout this paper all limits are taken as $n\to\infty$, allowing $p_{n}$ to diverge. Hereafter, we suppress the subscript $n$ when no confusion can arise. Write $\bm{\theta}_{0}=(\theta_{01},\ldots,\theta_{0p})^{\top}$. Define the index sets
@@
where $\theta_{\min}:=\min_{j\in T_{0}}|\theta_{0j}|$, $\gamma>0$ is a fixed constant independent of $n$ and $p$, and $\eta_{p}$ is a positive sequence depending on $p$. We write $s_{0}:=|T_{0}|$ for the support size, which may diverge, and $s_{\gamma}:=|T_{\gamma}|$ for the number of coefficients of moderate magnitude, kept fixed as $n$ varies.
@@
    There exists $\theta_{\max}>0$, independent of $n$ and $p$, such that $\|\bm{\theta}_{0}\|_{\infty}\leq\theta_{\max}$.
@@
Assumption 1-(ii) states that the non-zero coefficients outside $T_{\gamma}$ are precisely the weak signals, and Assumption 1-(iii) requires that their aggregate magnitude is asymptotically negligible. Under Assumption 1, $(\bm{\theta}_{0,n}^{\top},0,0,\ldots)\in\mathbb{R}^{\infty}$ lies in $\ell^{2}$ uniformly in $n$, because
@@
We define the ordinary least squares estimator $\tilde{\bm{\theta}}_{n}=(\tilde{\theta}_{n1},\ldots,\tilde{\theta}_{np})^{\top}$ after variable selection as the solution to the estimating equation
@@
Under the assumption that $s_{\gamma}$ is independent of $p$, the following lemma holds.
@@
Fujimori and Tsukuda (2026) analyze asymptotic normality for high‑dimensional estimators via weak convergence in $\ell^{2}$, including settings with a possibly infinite‑dimensional nuisance parameter and dependent observations. Their analysis is confined to the fixed-sparsity regime, which materially simplifies the arguments. In contrast, we allow $s_{0}\to\infty$ while keeping $s_{\gamma}$ fixed.
@@ 4 Proofs
$n^{-1}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\right\}^{2}\leq\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}\|_{2}^{2}\left\|n^{-1}\sum_{i=1}^{n}\bm{Z}_{i}\right\|_{2}^{2}=o_{p}(1)O_{p}(1)=o_{p}(1).$

### v3

@@ 1 Introduction
Throughout the paper, we use the following notation. We write $k]=\{1,\ldots,k\}$ for $k\in\mathbb{N}$. For $q\in\mathbb{N}\cup\{\infty\}$, write $\|\bm{v}\|_{q}$ for the $\ell^{q}$ norm: for $\bm{v}=(v_{1},\ldots,v_{p})^{\top}\in\mathbb{R}^{p}$ with $p\in\mathbb{N}$, $\|\bm{v}\|_{q}=(\sum_{j=1}^{p}|v_{j}|^{q})^{1/q}$ $(q<\infty)$, $\|\bm{v}\|_{\infty}=\max_{1\leq j\leq p}|v_{j}|$. For an infinite sequence $\bm{v}=(v_{j})_{j\geq 1}\in\ell^{q}$, the same formulae apply. For a bounded linear operator $\bm{A}:\ell^{2}\to\ell^{2}$ or a matrix $\bm{A}$ in finite dimensions, let $\|\bm{A}\|_{\mathrm{op}}=\sup_{\|\bm{x}\|_{2}=1}\|\bm{A}\bm{x}\|_{2}$ denote the $\ell^{2}$‑induced operator norm. For a vector $\bm{v}=(v_{j})_{j\geq 1}$ and an index set $T\subset\mathbb{N}$ with $|T|<\infty$, write $\bm{v}_{T}=(v_{j})_{j\in T}$, which is viewed as an element of $\mathbb{R}^{|T|}$. In the finite‑dimensional case $\bm{v}\in\mathbb{R}^{p}$, we restrict to $T\subset[p]$. Similarly, for a $p\times p$ matrix $\bm{A}=(A_{ij})_{i,j\in[p]}$ and index sets $T,T^{\prime}\subset[p]$, write $\bm{A}_{T,T^{\prime}}=(A_{ij})_{i\in T,j\in T^{\prime}}$. For a random variable $X$, we write the sub-Gaussian norm of $X$, if it exists, as $\|X\|_{\psi_{2}}$, that is, $\|X\|_{\psi_{2}}=\inf\{C>0:\,\mathsf{E}[\exp({X^{2}}/{C^{2}})]\leq 2\}$. For a $d$-dimensional random vector $\bm{X}$, the sub-Gaussian norm of $\bm{X}$ is $\|\bm{X}\|_{\psi_{2}}=\sup_{\|\bm{u}\|_{2}=1}\|\bm{u}^{\top}\bm{X}\|_{\psi_{2}}$. For $d\in\mathbb{N}$, $\bm{\mu}\in\mathbb{R}^{d}$, and $\bm{V}\in\mathbb{R}^{d\times d}$ positive semi‑definite, we write $\mathcal{N}_{d}(\bm{\mu},\bm{V})$ for the $d$-dimensional normal distribution with mean $\bm{\mu}$ and covariance $\bm{V}$, with the subscript omitted when $d=1$. For random elements $X$ and $Y$ and a sequence $\{X_{n}\}_{n\in\mathbb{N}}$ of random elements, we write $X\stackrel{{\scriptstyle d}}{{=}}Y$ if $X$ and $Y$ have identical distributions, and we write $X_{n}\to^{d}X$ and $X_{n}\to^{p}X$ for convergence in distribution and in probability as $n\to\infty$, respectively.
@@ 2 Model set‑up and main result
where $\bm{\theta}_{n}\in\mathbb{R}^{p_{n}}$ and $\bm{Z}_{i,n}=(Z_{i1,n},\ldots,Z_{ip_{n},n})^{\top}\in\mathbb{R}^{p_{n}}$ for $i\in[n]$. Assume that $\{(\bm{Z}_{i,n},\epsilon_{i,n})\}_{i\in[n]}$ are i.i.d. with $\mathsf{E}(\bm{Z}_{1,n})=\bm{0}$, $\mathsf{E}(\epsilon_{1,n})=0$, $\mathsf{V}(\epsilon_{1,n})=\sigma_{n}^{2}>0$, and $\bm{Z}_{1,n}$ independent of $\epsilon_{1,n}$. Let $\bm{\theta}_{0,n}$ denote the true parameter vector. Unless otherwise stated, throughout this paper all limits are taken as $n\to\infty$, allowing $p_{n}$ to diverge. Hereafter, we suppress the subscript $n$ when no confusion can arise. Write $\bm{\theta}_{0}=(\theta_{01},\ldots,\theta_{0p})^{\top}$. Define the index sets
@@
where $\theta_{\min}:=\min_{j\in T_{0}}|\theta_{0j}|$, $\gamma>0$ is a fixed constant independent of $n$ and $p$, and $\eta_{p}$ is a positive sequence depending on $p$. We write $s_{0}:=|T_{0}|$ for the support size, which may diverge, and $s_{\gamma}:=|T_{\gamma}|$ for the number of coefficients of moderate magnitude.
@@
    There exists $\theta_{\max}>0$, independent of $n$ and $p$, such that $\|\bm{\theta}_{0,n}\|_{\infty}\leq\theta_{\max}$.
@@
Assumption 1-(ii) states that the non-zero coefficients outside $T_{\gamma}$ are precisely the weak signals, and Assumption 1-(iii) requires that their aggregate magnitude is asymptotically negligible. Under Assumption 1, $(\bm{\theta}_{0,n}^{\top},0,0,\ldots)^{\top}\in\mathbb{R}^{\infty}$ lies in $\ell^{2}$ uniformly in $n$, because
@@
The post-selection ordinary least squares estimator $\tilde{\bm{\theta}}_{n}=(\tilde{\theta}_{n1},\ldots,\tilde{\theta}_{np})^{\top}$ is defined as the solution to the estimating equation
@@
Fujimori and Tsukuda (2026) analyze asymptotic normality for high‑dimensional estimators via weak convergence in $\ell^{2}$, including settings with a possibly infinite‑dimensional nuisance parameter and dependent observations. Their analysis is confined to the fixed-sparsity regime, which materially simplifies the arguments. In contrast, we allow $s_{0}\to\infty$ while keeping $T_{\gamma}$ fixed.
@@ 4 Proofs
$n^{-1}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\right\}^{2}\leq\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}\|_{2}^{2}n^{-1}\sum_{i=1}^{n}\|\bm{Z}_{iT_{\gamma}}\|_{2}^{2}=o_{p}(1)O_{p}(1)=o_{p}(1).$